OpenStudy (anonymous):

forctored, the expression 16x^2-25y^2 is equilavnt to

OpenStudy (therealmeeeee):

(4x - 5y)(4x+5y) <==ANSWER

OpenStudy (anonymous):

than you!

OpenStudy (anonymous):

@TheRealMeeeee yeah, u are right..

OpenStudy (therealmeeeee):

ha ty

OpenStudy (therealmeeeee):

well a medal would be nice for me help

OpenStudy (therealmeeeee):

ty

OpenStudy (therealmeeeee):

let me now if u need anymore help i will be gladly to help @alazaramos06

OpenStudy (anonymous):

@TheRealMeeeee jack bought 3 slice of cheese pizza and 4 slices of mushroom pizza for a total cost of $8.50. what is the cost of one slice of mushroom pizza?

OpenStudy (therealmeeeee):

on it

OpenStudy (therealmeeeee):

this is confusing do the question separably for now on ok

OpenStudy (therealmeeeee):

3c + 4m = 12.50 3c + 2m = 8.50 3c + 4m = 12.50 -3c -2m = -8.50 ------------------ 2m = 4.00 m = 2.00

OpenStudy (anonymous):

what is the product o -3x^2y and (5xy^2 + xy)?

OpenStudy (anonymous):

@TheRealMeeeee ^^^

OpenStudy (therealmeeeee):

wat

OpenStudy (therealmeeeee):

did u read my answer

OpenStudy (anonymous):

@TheRealMeeeee the q above

OpenStudy (therealmeeeee):

Simple Algebra answer is: 5xy/(5x^2-10xy) + 8xy/(2y^2-xy) = 5xy/ 5x(x-2y) + 8xy/y(2y-x) = y/ (x-2y) + 8x/(2y-x) = y/ (x-2y) - 8x/(x-2y) = ***(y-8 x)/(x-2 y)*** The Derivative: ((partial d))/((partial d)x)((5 x y)/(5 x^2-10 x y)+(8 x y)/(2 y^2-x y)) = (5 y)/(5 x^2-10 x y)-(5 x y (10 x-10 y))/(5 x^2-10 x y)^2+(8 x y^2)/(2 y^2-x y)^2+(8 y)/(2 y^2-x y) The Indefinite Integral: integral ((5 x y)/(5 x^2-10 x y)+(8 x y)/(2 y^2-x y)) dx = -15 y log(x-2 y)-8 x+16 y+constant

OpenStudy (therealmeeeee):

5xy/(5x^2-10xy) + 8xy/(2y^2-xy) =5xy/ 5x(x-2y) + 8xy/y(2y-x) =y/ (x-2y) + 8x/(2y-x) =y/ (x-2y) - 8x/(x-2y) =(y-8x)/(x-2y)